Clock

Version 2.0 (SHAREWARE)

Update of Clock V1.40 Released: 29.03.1993

Copyright 1992,1993 by Bernd Grunwald Zugspitzstr.78 W-8031 Eichenau Germany (Europe)

The passing-on will be only allowed if

the author’s name and the note of the copyright have NOT been removed,

if the passing-on is free or ONLY the copy-fees are covered and this doc-file is always included !

The passing-on of the Key-File is NOT allowed! THE AUTHOR WILL NOT BE LIABLE FOR

ANY DAMAGES CAUSED BY THIS PROGRAM

Contents

1. What is Clock ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2. Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3. How to start the program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

4. Project menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4.1. Date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4.2. Memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4.3. Select Alert Prefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.4. Other Prefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.5. Snapshot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4.5.1. For Kick 1.3 user !! (WARNING) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4.6. Save Config File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4.7. Set TaskPriority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4.8. Change Colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4.9. Info ... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.10. Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

5. Registration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

6. History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

IMPORTANT FOR THOSE WHO ALREADY USE THE CLOCK-PROGRAM

Please delete the config-file in the S: -directory since the size has changed. (Sorry, but I did not include spaces, ’cos I thought that the config-file should always have that size which is really necessary.)

1. What is Clock ?

Clock is a simple Clock-prog (what else :-) which can be started from the CLI as well as from the Workbench. It’s not extraordinary but it includes a little feature: Clock is able to pop up on the momentary screen automatically.

2. Installation

You can copy the icon wherever you want. The best place will be your boot-disk. The program can be started from the Workbench as well as from the CLI. Now you can also copy the ”ClockConfig”-File on your S: -directory, ”sam” on your C: -directory and ”Glocke” (german word for Bell) on your L: - directory (only necessary if you want to use the enclosed ”Glocken”-module). Now you are perfectly safe in starting the program. Another possibility to install Clock is to double-click the Install Icon. If the directory ”SYS:WBStartup” exists Clock will be put there; otherwise it will be put in the c: directory. If you’ve got the registered Version of Clock your Keyfile will be put on your boot disc.

3. How to start the program

Workbench: simply double-click on the icon. You can also put Clock in your WBStartup drawer. CLI: you have the possibility to start the clock with the following parameters

”?” : a short description of the program

”-cConfig-File” : Any Config-File (70 chars max.) e.g. -cs:ClockConfig

”-sSreenName” : Name of the Screen, you like Clock to open on; Use SPACES without ” e.g. -sTHE ANSWER

4. Project menu

4.1. Date

Switching on this menu-item the clock-window widen and the actual date will appear beside the time. This option is ”saveable”.

4.2. Memory

If you choose this menu-item the clock-window widen again and you will get information about the free memory, subdivided to chip, fast and total-memory (C: F: T:) shown in bytes (or kilobytes on lores screens). You can save this option too.

4.3. Select Alert Prefs

Here appears a window that looks like that:

Let’s start with the Gadget on the top left: If this Gadget is activated that means that the alert number 1 is active. In the next Gadget (|**:**:**|) you can enter the alert time. The time is needed in this format:

HH:MM:SS (Hours Minutes Seconds) (24h)

e.g.: 16:20:59 means: 4:20 PM and 59 seconds !

If you’ve got the reqtools.library in your LIBS: directory you can use the next Gadget. It will appear a file-requester where you can choose a program that should be started if the alert time is reached. This name will be put in the next String Gadget. If you’ve got a nice Sound-program (e.g. sam, that is included in this packet \begin{matrix}c \\ ◯ \end{matrix} in 1988 by Nic Wilson) that you like to use as bell just type the name in the string-gadget. The length of a filename inclusively its parameter is restricted to 70 chars. If you want to start a program while the clock shall go on running, you have to choose the next (”r”) gadget. It’s necessary that the file RUNBACK exists in your directory C: (If you don’t have RUNBACK you can also use RUN but first you have to rename a copy of RUN to RUNBACK). By using the sound-gadget (”s”) you can choose if you want to have the sinus sound as alarm or not.

The next line includes 2 Gadgets: the first from left says if a Requester should pop up, when the alert-time is reached. The Text of that Requester you choose by typing it in the string-gadget that is in the right of the first gadget in this line.

The next lines are for alert number 2 - 4.

If you want the changes to be used click ”Go On”, else click ”Cancel”.

4.4. Other Prefs

Here appears a window that looks like that:

Let’s start with the Gadget on the top left: If this Gadget is activated that means that the Alert on the beginning of every hour is active. If you’ve got the reqtools.library in your LIBS: directory you can use the next Gadget. It will appear a file-requester where you can choose a program that should be started if the alert time is reached. This name will be put in the next String Gadget. If you’ve got a nice Sound-program (e.g. sam, that is included in this packet \begin{matrix}c \\ ◯ \end{matrix} in 1988 by Nic Wilson) that you like to use as bell just type the name in the string-gadget. The length of a filename inclusively its parameter is restricted to 70 chars. The RUNBACK Option is activated and the Sound Option is not activated.

The next line includes 2 Gadgets: the first from left says if a Requester should pop up, when the alert-time is reached. The Text of that Requester you choose by typing it in the string-gadget that is in the right of the first gadget in this line.

The last line (under Name of Screen) includes 2 Gadgets: the first from left says if Clock should pop up on a special Screen. In the next string-gadget you can enter the Name of the screen where Clock should pop up. This option is case sensitive! Only the chars entered in the string-gadget are compared to the screen name. That means if you’ve got a screen with the name ”THE ANSWER 2.03 - DER ..” you only have to type ”THE ANSWER” or ”THE” ... .

If you want the changes to be used click ”Go On”, else click ”Cancel”.

4.5. Snapshot

If you want to fix the window on the momentary active screen you have to choose ”Snapshot” in the projectmenu or press the left Amiga-key + f. When fixing is activated you can see a check-mark in front of the menu-item. After a second use of this menu-item the check-mark disappears again and the window is not fixed on the screen anymore. If you have not activated the snapshot-function and an additional window is opened beside the clock-window, both will appear on the momentary screen you have changed to.

If you have chosen ”Snapshot” you can click the window onto the next screen with the help of the gadget in the clock window. To make it easier I have also introduced a key- combination (CTRL + left-ALT) which helps you to bring the clock-window in the fixed form on your momentary screen (press both keys for about a second).

4.5.1. For Kick 1.3 user !! (WARNING)

If you use Kick1.3 and want to fix the clock-window on a screen, you are not allowed to close it without having clicked the window with the help of the gadgets on the next screen. If you don’t do this you will never see a clock window again and probably will have a guru ! (unfortunately you could only abolish these failure by letting clock run just one time. But I believe it’s more important to have several clocks at the same time and most persons who use clock also use Kick 2.0 or greater as I do.) PS. I haven’t tested Clock under 1.3, because I now use Kick 2.0 and I don’t know if it works correctly

under 1.3. Because I have not used any Kick 2.0 only Library function I hope it still works under Kick 1.3.

4.6. Save Config File

Of course you can put the window of the clock to a place you like. When you find a position you like you can save it with the menu-item ”save config”. This means that the program will appear on this place every time you start Clock. Besides you can save the following items: Priority, Sound on/off, alarm-file, fixing on/off, date on/off, memory on/off, screen name and the alarm-times (!) This produces a 868 byte file called ”ClockConfig” in the S: directory. If you’ve got the reqtools.library in your LIBS: directory it appears a requester, where you can choose the Place and name of the config file. (see also ”How to start the program”)

4.7. Set TaskPriority

If you think the clock takes too much CPU-time you are able to change the task-priority. Just choose the ”Set TaskPriority”-Item from the menu and you will get a window where you can enter a new priority. The range is from -9 to 9 where greater values than 9 will be truncated and only the first digit will be used.

4.8. Change Colors

If you have the reqtools.library in your LIBS: directory you can change the colors of the screen the clock window is opened on.

4.9. Info ...

One of the most important menu-items. :-) It just shows my address...

4.10. Quit

Well, this is an undocumented feature of Clock ! :-)

5. Registration

If you use Clock longer than 1 week you have to register. Then you will either get a Key-File via Fido or Z-Net, or a disc with the new Clock Version, key-file and a few little bonus programs. If you’ve got the key-file the Info-Window will not be shown anymore in the beginning. The key-file will also be usable (who knows; see Keyfilemaker for LHA ... ) with later versions of Clock so I think it is affordable for everyone. If you then get a new Clock Version via Net or PD disc you can use it without any windows or requesters going on your nerves. In the key-file your Name and address are saved, so that I can quickly say from whom this file is. So don’t pass-on this key-file, because it is NOT allowed!

PS. The newest Clock Version will be put in the ”Amiga City München” Mailbox where you can easily get it from like this: As Username you type ”Gast” (german word for guest) then you type ”BR SAUGER/AMIGA” and with ”INH **” you can take a look at the files that are in that area. To download the newest Clock Version with Zmodem you just type ”LE xx” (where xx = is the number of Clock in the filelist).

The number of the AMC: 1200 - 14400 Bps HST V32 V32bis V42 V42bis Tel: +49-89-7696801 24h Online 1200 - 2400 Bps MNP2-5 V42 V42bis Tel: +49-89-7608567 24h Online

For the costs of registration see the included registration form. (please only use this form)

IF THE PROGRAM DOESN’T WORK AS ITS NAME PROMISES, PLEASE SEND ME YOUR EQUIPMENT-LIST (COMPUTER, KICK-VERSION, WORKBENCH, STARTUP-SEQUENCE) !

If you have any suggestions, bug-reports, ideas etc. you can mail me via: Zerberus-Net : EMBO@AMC.ZER Fidonet : Bernd Grunwald of 2:246/221.100

Have fun and enjoy this little proggy !

Many thanx to:Holgi (HOLGI@CORONA.ZER)

Peter Fröhlich (P.FROEHLICH@AMC.ZER)

Thomas Strauß (THOMAS STRAUSS@SAARAG.ZER) Stefan Brandes Ralf for translating this document (better late than never :-) If there are errors in this doc don’t blame Ralf for that because I changed much of that doc ... and all others people, computers, coffee-cups...

6. History

Changes since V 1.40: -Locale library is used if possible (who wants to translate the texts in other languages than german and english?) -4 Alerttimes saveable \underline{6. History 7}

-requester can pop up on alert -clock is able to pop up on a screen with given screen-name -Hourly Signal -clock now is Shareware ...

Changes since V 1.35k: -color change in ”set alarm prefs” changed in a separate menu-item -show date in the window-title added -show memory in the window title added -sound-gadget in ”Alarm Preferences” added -runback-gadget in ”Alarm Preferences” added -string-gadget will now activated if necessary (e.g. set alarm-time) -for 2.0-user: now the menus have the correct font -alarm-time is ”saveable” (on request by one user :-) -for 1.3-user: window-size corrected -some minor bugs and news again...

Changes since V 1.35h: -Autodetach added -CPU-usage reduced by 50 percent :-) -file-requester for alarmfile-selection added -color-requester added -date- and memory-display added -bug in saving the window-position corrected (hope so :-) -parameter ”sound on/off” and ”alarm-file” added -ctrl and left-alt pops up the window on the momentary screen (hold down for a second) -some minor bugs and news (to be continued...)

Changes since V 1.34d: -now Kick 2.0-compatible, I think ;-)

-any program may be started on alarm -now you can switch off the alarm-sound -no ”ghosting” window under kick1.3 -automatically uses the right size for the menu-fonts if you are not using the systemfonts -some minor bugs removed (kick1.3) -gadget for ”klicking-on-the-next-screen” -bug removed: there were screens less than the x- and y- coordinates from clock. In these cases clock makes a journey to India :-(

that’s all, folks !